OpenStudy (anonymous):

Which equation does the graph of the systems of equations solve? *2 linear graphs. They intersect at 0,3 Answer choices: (a) -2x + 3 = 8x - 3 (b) -2x + 4 = -8x + 4 (c) 2x + 4 = 8x - 4 (d) 2x + 3 = -8x + 3

OpenStudy (anonymous):

D. 2x+3 = -8x+3

OpenStudy (mathmale):

You are saying that the system of equations intersect at (0,3). (Please remember to use those parentheses.) One way to respond would be to test each of the four given possible solutions by substituting x=0 and y=3. For example, I'm choosing the fourth possible answer, 2x + 3 = -8x + 3. I let x = 0 and determine whether the left and right sides of this equation are indeed equal, as well as = to 3 (that is, y=3). Outcome: 2(0)+3 = =8(0) +3 => the left side is 3 and the right side is also 3. It just so happened that I lucked out in choosing the fourth option. (0,3) is indeed a solution of that system of linear equations. Note that even though it is 2x + 3 = -8x + 3 that is given, you can revert to the "system of linear equations" by letting each side = y: y = 2x + 3 and y=-8x + 3. You could try solving this system of linear equations yourself to determine that the solution truly is (0,3).

OpenStudy (anonymous):

Thank you both! Very helpful.